# Priors from Envisioned Posterior Judgments: A Novel Elicitation Approach With Application to Bayesian Clinical Trials

### v1

@@ 2 Methods
Formally, let $\theta$ denote the parameter(s) of interest (e.g., the mean difference in outcomes between intervention and control arms), $f(\theta;a)$ be a family of prior distributions for $\theta$ indexed by a set of fixed parameter values $a$, $f(Y\mid\theta)$ be the sampling model/likelihood function, and $f(\theta\mid Y)\propto f(Y\mid\theta)f(\theta)$ be the posterior distribution for $\theta$ given $Y$. Our proposed elicitation method consists of two steps:
@@
In the case study presented below, the goal was to elicit a prior for a (univariate) mean parameter $\theta$ from a Normal data model, i.e., $Y\mid\theta\sim\mathcal{N}\left(\theta,1/\tau\right)$, where $\tau$ is the precision parameter. For the $k$-th scenario, experts were asked to provide their response $\tilde{\theta_{k}}$ given data $\mathbf{Y}_{k}$, comprised of a sample of $n_{k}$ observations for $Y$. We assumed a conjugate Normal prior $\mathcal{N}\left(\mu_{0},1/\tau_{0}\right)$ for $\theta$, and our elicitation goal was to infer the value of $a\equiv(\mu_{0},\tau_{0})$. Under these assumptions, the posterior mean $\mu_{k}=(\tau_{0}\mu_{0}+n_{k}\bar{y}/s^{2})/(\tau_{0}+n_{k}/s^{2})$, where $\bar{y}$ is the sample mean and $s^{2}$ is the sample variance. With a squared difference as the discrepancy function, the inferred value for ($\mu_{0}$, $\tau_{0}$) was that value that minimizes the RMSD. This optimization was straightforward as it involved only two variables and a well-behaved objective function.
@@ 3.4.3 Initial elicitation data collection and analysis
Experts were asked to complete the questionnaire by providing their best estimate for the mean change in 6WMT from 2021 to 2022 in the UC group (i.e., $\ tilde{theta_{k}}$) if they were to observe the outcomes from each of the 16 hypothetical scenarios. Our statisticians did not directly communicate with the experts during this stage; instead, the PI handled all communication. After receiving the elicited responses, we inferred the prior distribution for each expert and assessed their level of coherency using the method described in section 2.3.
@@ 4.1 Initial elicitation
Three experts (1, 2, 3) achieved a RMSD less than 3 metres. Note, though, that the expert who achieved the highest coherency (1) had an inferred prior mean of -0.3 and SD of 0.2, indicating they were nearly certain that the mean change was very close to zero. For this expert, a concern would be whether this degree of certainty could be justified, despite the judgments being coherent. Examination of the individual responses for expert 10 strongly suggested the expert misunderstood what was requested, which would explain their large values for the prior SD and the incoherency measures.

### v2

@@ 2 Methods
Formally, let $\theta$ denote the parameter(s) of interest (e.g., the mean difference in outcomes between intervention and control arms), $f(\theta;a)$ be a family of prior distributions for $\theta$ indexed by a set of fixed parameter values $a$, $f(Y\mid\theta)$ be the sampling model/likelihood function, and $f(\theta\mid Y;a)\propto f(Y\mid\theta)f(\theta;a)$ be the posterior distribution for $\theta$ given $Y$. Our proposed elicitation method consists of two steps:
@@
In the case study presented below, the goal was to elicit a prior for a (univariate) mean parameter $\theta$ from a Normal data model, i.e., $Y\mid\theta\sim\mathcal{N}\left(\theta,1/\tau\right)$, where $\tau$ is the precision parameter. For the $k$-th scenario, experts were asked to provide their response $\tilde{\theta_{k}}$ given data $\mathbf{Y}_{k}$, comprised of a sample of $n_{k}$ observations for $Y$. We assumed a conjugate Normal prior $\mathcal{N}\left(\mu_{0},1/\tau_{0}\right)$ for $\theta$, and our elicitation goal was to infer the value of $a\equiv(\mu_{0},\tau_{0})$. Under these assumptions, the posterior mean is $\mu_{k}(a)=(\tau_{0}\mu_{0}+n_{k}\bar{y}/s^{2})/(\tau_{0}+n_{k}/s^{2})$, where $\bar{y}$ is the sample mean and $s^{2}$ is the sample variance. With a squared difference as the discrepancy function, the inferred value for ($\mu_{0}$, $\tau_{0}$) from minimizing the RMSD is
$(\hat{\mu_{0}},\hat{\tau_{0}})=\arg\min_{\mu_{0},\tau_{0}}\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left(\tilde{\theta_{k}}-(\tau_{0}\mu_{0}+n_{k}\bar{y}/s^{2})/(\tau_{0}+n_{k}/s^{2})\right)^{2}}$
This optimization is straightforward as it involves only two variables and a well-behaved objective function.
@@ 3.4.3 Initial elicitation data collection and analysis
Experts were asked to complete the questionnaire by providing their best estimate for the mean change in 6WMT from 2021 to 2022 in the UC group (i.e., $\tilde{\theta_{k}}$) if they were to observe the outcomes from each of the 16 hypothetical scenarios. Our statisticians did not directly communicate with the experts during this stage; instead, the PI handled all communication. After receiving the elicited responses, we inferred the prior distribution for each expert and assessed their level of coherency using the method described in section 2.3.
@@ 4.1 Initial elicitation
Three experts (1, 2, 3) achieved a RMSD less than 3m. Note, though, that the expert who achieved the highest coherency (1) had an inferred prior mean of $-0.3$m and SD of 0.2m, indicating they were nearly certain that the mean change was very close to zero. For this expert, a concern would be whether this degree of certainty could be justified, despite the judgments being coherent. Examination of the individual responses for expert 10 strongly suggested the expert misunderstood what was requested, which would explain their large values for the prior SD and the incoherency measures.